OpenStudy (anonymous):

Given the set {1, 2, 3, 4, 5} how many different subsets with three numbers in each can I make?

OpenStudy (anonymous):

ooh subsets of 3 numbers you are being asked how many ways you can choose 3 items from a set of 5 this is called "5 choose 3" written as \(\binom{5}{3}\) or \(_5C_3\) and computed via \[\binom{5}{3}=\frac{5\times 4}{2}\]

OpenStudy (anonymous):

Thank you! It could also be written as \[\frac{5 * 4 * 3}{3 * 2 *1}\], right?